# Mamba Neural Operator: Who Wins? Transformers vs. State-Space Models for PDEs

### v2

@@ Abstract
Partial differential equations (PDEs) are widely used to model complex physical systems, but solving them efficiently remains a significant challenge. Recently, Transformers have emerged as the preferred architecture for PDEs due to their ability to capture intricate dependencies. However, they struggle with representing continuous dynamics and long-range interactions. To overcome these limitations, we introduce the Mamba Neural Operator (MNO), a novel framework that enhances neural operator-based techniques for solving PDEs. MNO establishes a formal theoretical connection between structured state-space models (SSMs) and neural operators, offering a unified structure that can adapt to diverse architectures, including Transformer-based models. By leveraging the structured design of SSMs, MNO captures long-range dependencies and continuous dynamics more effectively than traditional Transformers. Through extensive analysis, we show that MNO significantly boosts the expressive power and accuracy of neural operators, making it not just a complement but a superior framework for PDE-related tasks, bridging the gap between efficient representation and accurate solution approximation.
I Introduction
@@
Partial differential equations (PDEs) describe various real-world phenomena, such as heat transfer (Heat Equation), fluid dynamics (Navier-Stokes), and biological systems (Reaction-Diffusion). While analytical solutions are sought, many PDEs—like the Navier-Stokes equations—lack closed-form solutions, making them computationally intensive to solve. Numerical methods, such as finite element, finite difference [1], and spectral methods, discretise these equations but involve trade-offs between computational cost and accuracy. Coarser grids reduce computational load but sacrifice precision, while finer grids increase both accuracy and computational expense. Recent advancements in deep learning have changed techniques for solving PDEs. Physics-Informed Neural Networks (PINNs) [2, 3] integrate governing equations and boundary conditions into the loss function, but often struggle with generalisation and require retraining for changes in coefficients. Neural operators[4, 5], on the other hand, learn mappings between function spaces, offering a mesh-free, data-driven approach that generalises better across different PDE instances.
@@
Operator learning has gained traction with models like DeepONet[6] and the Fourier Neural Operator (FNO)[7], which achieved state-of-the-art performance. These models learn input-output mappings to approximate complex operators, similar to sequence-to-sequence problems. Transformers[8] have become a go-to architecture for PDEs[9, 10, 11] due to their ability to capture long-range dependencies. However, their quadratic complexity limits efficiency for tasks such as long-time integration. To overcome this, efficient variants like Galerkin attention[9] reduce computational cost to linear scaling. While these models improve efficiency, they trade off model capacity by approximating the self-attention mechanism, potentially reducing accuracy for tasks that need precise attention. Moreover, Transformers face challenges with PDEs due to limited context windows, inefficiency with continuous data, and high memory usage, making them less effective for capturing dependencies over continuous domains and high-resolution grids.
@@
While Transformers are popular for PDE modelling, they have limitations in handling continuous data and high-resolution grids. An emerging alternative is State-Space Models (SSMs)[12, 13], which offer better scalability, reduced memory usage, and improved handling of long-range dependencies in continuous domains compared to Transformers. In particular, Mamba[14] is a novel way designed to effectively capture long-range dependencies, handle continuous data efficiently, and reduce memory consumption in sequence-to-sequence problems. Although Transformers dominate applications like foundational models and computer vision, the use of SSMs—especially Mamba—for neural operators in PDEs remains underexplored, and their theoretical connections and potential advantages are yet to be fully understood.
@@ II Related Work
Data-Driven PDEs. Recent advances in fluid dynamics and solving PDEs have led to architectures modelling continuous-time solutions and multiparticle dynamics[15, 16]. Physics-informed models now offer solutions in unsupervised and semi-supervised settings[2, 7]. These models typically encode spatial data and evolve over time, utilising methods like convolutional layers[17, 18] symbolic neural networks[19], and residual networks[20]. Finite element methods (FEM), including Galerkin and Ritz, are also integrated into learning frameworks[21].
@@
Neural Operators. Neural operators, such as the Graph Neural Operator[22] and Fourier Neural Operator[7], excel at learning mappings in infinite-dimensional spaces, particularly by leveraging techniques like graph structures or transformations in Fourier space. The Fourier Neural Operator (FNO) and its variations, including the incremental, factorised, adaptive FNO, and FNO+[23, 24, 25] have shown exceptional performance in both speed and accuracy. Their key advantage lies in their ability to maintain discretisation invariance, which sets them apart in many applications. DeepONet[6] pioneered the nonlinear operator approximation using separate networks for inputs and query points, while extensions like MIONet handle multiple inputs[26]. Challenges like irregular grids are being addressed through grid mapping and subdomain partitioning[27, 28] though scalability for diverse inputs remains a key focus.
@@
Transfomers for PDEs. The Transformer model[8] stands out due to its distinctive features, primarily its use of attention mechanisms to model the relationships among input elements. Initially, it was developed for NLP, and attention mechanisms have been adapted to PDEs, providing flexible and efficient mappings between function spaces. Galerkin attention[9] introduced linear complexity to reduce computational costs, inspiring further developments like GNOT[29] and OFormer[10], which achieve state-of-the-art results. Additionally, graph-based Transformers have also been explored to capture complex interactions in irregular domains[11].
@@
State-Space Models for PDEs & Comparison to Ours. Initial studies on SSMs for PDEs, like MemNO[30], explored combining FNO with S4 but were restricted to low-resolution or noisy inputs. In contrast, we introduce the Mamba Neural Operator, which generalises the SSM framework to neural operators, making it compatible with any architecture, including Transformers. Our approach extends the theoretical foundations for broad applicability to any PDE family, highlighting Mamba’s effectiveness in diverse scenarios. At the time of our submission, the work of that[31] proposed integrating state-space models into neural operators for dynamical systems. While related, our work differs significantly, their approach focuses on dynamical systems and tests only on ordinary differential equations (ODEs), whereas we target parametric partial differential equations (PDEs). Additionally, we provide a theoretical understanding showing that neural operator layers share a comparable structural framework with time-varying SSMs, demonstrating alignment between hidden space updates and the iterative process in neural operators.
III Mamba Neural Operator
@@ III-A Problem Statement
III-B Preliminaries: Transformer and Mamba
@@
of which the memory complexity is $O(n^{2})$. To reduce the computational inefficiency, Galerkin-type attention was proposed by[9] to remove Softmax attention with linear complexity. It defines as follows:
$\operatorname{Attention}_{g}(Q,K,V)=\frac{Q(\widetilde{K}^{T}\widetilde{V})}{d},$ (4)
in which $\widetilde{\cdot}$ denotes layer normalisation, as described in[32]. The Galerkin-type attention mechanism involves two matrix product operations, resulting in a computational complexity of $O(nd^{2})$. This reduces the sequence length dependency to only $O(n)$.
@@
where $A\in\mathbb{C}^{N\times N}$ and $B,C\in\mathbb{C}^{N}$. Mamba, a more advanced variant of SSMs, refines this formulation by incorporating efficient state space parameterisation and selection mechanisms. Unlike earlier models such as S4, which uses bilinear method, Mamba adopts zero-order holds, allowing it to handle larger hidden states and longer sequences more effectively. This makes Mamba particularly well-suited for complex sequence modeling tasks, such as natural language processing and time-series analysis.
III-C State Space Models Discretisation for PDEs
@@
We start by demonstrating that the discretisation of S6 (Mamba) is equivalent to the well-known Eular method when the Taylor series expansion is applied. Mamba utilises zero-order holds, resulting in the following discretisation, which reads:
@@
Discretisation of SSM. To incorporate the SSM into deep learning frameworks, we need to transform the continuous-time SSM into a discrete formulation. This is done by expressing the continuous-time system as an ordinary differential equation (ODE) and then solving it numerically. As discussed in[33], the discrete SSM reads:
@@
$\begin{split}h(t+\Delta t)&=\hat{A}h(t)+\hat{B}u(t)=(I+\Delta A)h(t)+\Delta Bu(t)\\ &=h(t)+\Delta(Ah(t)+Bu(t))=h(t)+\Delta h^{\prime}(t).\end{split}$ (9)
which implies it is a first order eular method. It is straightforward to show that $\tilde{A}=\hat{A}$ since $\tilde{A}=e^{A\Delta}=I+A\Delta+O(\Delta^{2})=I+A\Delta=I+A\Delta=\hat{A}$. Similarly, we observe that $\tilde{B}=\bar{B}=\hat{B}$. Therefore, the discretisation used in the SSM method can be replaced with the zero-order hold method by substituting $\hat{A}=\tilde{A}$ and $\hat{B}=\tilde{B}$, we get:
$\begin{split}h(t+\Delta t)&=\hat{A}h(t)+\hat{B}u(t)=\tilde{A}h(t)+\tilde{B}u(t)\\ &=(e^{\Delta A})h(t)+((\Delta A)^{-1}(e^{\Delta A}-I)\cdot\Delta B)u(t)\\ &=(I+\Delta A+O(\Delta^{2})h(t)+(\Delta B)u(t)\\ &=(I+\Delta A)h(t)+(\Delta B)u(t)\\ &=h(t)+\Delta Ah(t)+\Delta Bu(t)\\ &=h(t)+\Delta(Ah(t)+Bu(t))=h(t)+\Delta h^{\prime}(t).\end{split}$ (10)
@@
Why is Proposition 1 important for PDEs? Proposition 1, which establishes the equivalence between the Zero-Order Hold (ZOH) method and the Euler method, is crucial for understanding Mamba’s performance in solving partial differential equations (PDEs). This equivalence demonstrates that ZOH can be viewed as a more generalised and accurate variant of the Euler method. Specifically, while the Euler method is derived by applying the Taylor series expansion and truncating it at the first order, the ZOH method retains additional higher-order terms from the Taylor series, making it inherently more accurate. This distinction has significant implications when solving PDEs. Higher-order methods like ZOH provide better approximations of a system’s behaviour without requiring excessively small step sizes $\Delta$, which are often necessary for the Euler method to achieve a similar level of accuracy. Smaller step sizes can result in increased computational cost and potential numerical instability. By utilising ZOH’s higher-order accuracy, the Mamba architecture can handle a wide range of step sizes, ensuring both stability and convergence without compromising precision.
III-D Network Architecture
@@
As depicted in Figure 1, the data processing pipeline in our Mamba Neural Operator (MNO) is composed of three key stages: Bi-Directional Scan Expand, S6/Cross S6 Block, and Bi-Directional Scan Merge. When solving PDEs over a fixed grid, the input data can be structured as grid-based data, similar to an image. In the first stage, Bi-Directional Scan Expand, the MNO unfolds the input data into sequences by traversing the grid along two distinct paths. These sequences, representing input patches, are processed independently in the next step. The second stage, S6/Cross S6 Block, involves processing each patch sequence using either an S6 or Cross S6 block, depending on the model variation being employed. For instance, in the enhanced version of Mamba, the GNOT model utilises a Cross S6 block followed by an S6 block for further refinement. Finally, in the Bi-Directional Scan Merge stage, the processed sequences are reshaped and merged back together to generate the output map, completing the data forwarding process. This structured approach allows the MNO to efficiently handle grid-based input data, enabling scalable solutions for PDEs.
@@
Next, the parameters $(\tilde{B},\tilde{C},\tilde{\Delta})$ are computed by combining the updated values from both inputs according to the following equations:
@@
III-E Mamba for Neural Operators
Neural Operators[22] aim to learn mappings between function spaces, providing a framework for solving partial differential equations (PDEs) and other problems involving continuous functions. It updates the value by an iterative method: $i_{0}\rightarrow i_{1}\rightarrow\ldots\rightarrow i_{T},$ where each $i_{j}$ (for $j=0,1,...,T-1$) maps to $\mathbb{R}^{d_{v}}$. Let the input be $a(x)$ and the output be $u(x)$ The input $a$, drawn from set $A$, is initially lifted to a higher-dimensional representation: $v_{0}(x)=P(a(x))$ where $P$ is a local transformation, typically parameterised by a fully-connected neural network. We then apply iterations to update $i_{t}\rightarrow i_{t+1}$ as defined in Definition 1. The final output: $u(x)=Q(v_{T}(x))$ is the result of projecting $v_{T}$ via the transformation: $Q:\mathbb{R}^{d_{v}}\to\mathbb{R}^{d_{u}}.$ Each update from $i_{t}$ to $i_{t+1}$ involves the integration of a non-local integral operator $K$ and a local nonlinear activation function $\sigma$. One of the main results of this work is establishing the equivalence between neural operators and the Mamba framework. Therefore, we first introduce fundamental definitions stated in [22] that are essential for demonstrating this relationship.
Figure 1: (A) Illustration of Mamba Neural Operator. Input image patches are processed by following two distinct scanning paths (referred to as Bidirectional -Scan). Each sequence generated from these paths is passed through separate S6 blocks/ Cross S6 Blocks for independent processing. Afterwards, the outputs from the S6 blocks / Cross S6 Blocks are combined to form a feature map, resulting in the final output (Bidirectional-Merge). (B) and (C) are the detailed block of the S6 Block and Cross S6 Block respectively. The detail network architecture and definition of Cross S6 Block can be found in Appendix A.

### v3

@@ Abstract
Partial differential equations (PDEs) are widely used to model complex physical systems, but solving them efficiently remains a significant challenge. Recently, Transformers have emerged as the preferred architecture for PDEs due to their ability to capture intricate dependencies. However, they struggle with representing continuous dynamics and long-range interactions. To overcome these limitations, we introduce the Mamba Neural Operator (MNO), a novel framework that enhances neural operator-based techniques for solving PDEs. MNO establishes a formal theoretical connection between structured state-space models (SSMs) and neural operators, offering a unified structure that can adapt to diverse architectures, including Transformer-based models. By leveraging the structured design of SSMs, MNO captures long-range dependencies and continuous dynamics more effectively than traditional Transformers. Through extensive analysis, we show that MNO significantly boosts the expressive power and accuracy of neural operators, making it not just a complement but a superior framework for PDE-related tasks, bridging the gap between efficient representation and accurate solution approximation. Our code is available on https://github.com/Math-ML-X/Mamba-Neural-Operator
keywords:
[…]
1 Introduction
@@
Partial differential equations (PDEs) describe various real-world phenomena, such as heat transfer (Heat Equation), fluid dynamics (Navier-Stokes), and biological systems (Reaction-Diffusion). While analytical solutions are sought, many PDEs—like the Navier-Stokes equations—lack closed-form solutions, making them computationally intensive to solve. Numerical methods, such as finite element, finite difference [mehra2010comparison], and spectral methods, discretise these equations but involve trade-offs between computational cost and accuracy. Coarser grids reduce computational load but sacrifice precision, while finer grids increase both accuracy and computational expense. Recent advancements in deep learning have changed techniques for solving PDEs. Physics-Informed Neural Networks (PINNs) [raissi2019physics, mattey2021physics] integrate governing equations and boundary conditions into the loss function, but often struggle with generalisation and require retraining for changes in coefficients. Neural operators[bhattacharya2021model, kovachki2023neural], on the other hand, learn mappings between function spaces, offering a mesh-free, data-driven approach that generalises better across different PDE instances.
@@
Operator learning has gained traction with models like DeepONet[lu2019deeponet] and the Fourier Neural Operator (FNO)[li2020fourier], which achieved state-of-the-art performance. These models learn input-output mappings to approximate complex operators, similar to sequence-to-sequence problems. Transformers[vaswani2017attention] have become a go-to architecture for PDEs[cao2021choose, li2022transformer, bryutkin2024hamlet] due to their ability to capture long-range dependencies. This is because the transformers address this challenge by providing a global receptive field: any two spatial tokens can interact in a single layer, making long-range effects as accessible as local ones. The attention mechanism further allows dynamic weighting, where the model adaptively selects the most relevant interactions regardless of distance. However, their quadratic complexity limits efficiency for tasks such as long-time integration. To overcome this, efficient variants like Galerkin attention[cao2021choose] reduce computational cost to linear scaling. While these models improve efficiency, they trade off model capacity by approximating the self-attention mechanism, potentially reducing accuracy for tasks that need precise attention. Moreover, Transformers face challenges with PDEs due to limited context windows, inefficiency with continuous data, and high memory usage, making them less effective for capturing dependencies over continuous domains and high-resolution grids.
@@
While Transformers are popular for PDE modelling, they have limitations in handling continuous data and high-resolution grids. An emerging alternative is State-Space Models (SSMs)[gu2021combining, gu2022parameterization], which offer better scalability, reduced memory usage, and improved handling of long-range dependencies in continuous domains compared to Transformers. In particular, Mamba[gu2023mamba] is a novel way designed to effectively capture long-range dependencies, handle continuous data efficiently, and reduce memory consumption in sequence-to-sequence problems. Although Transformers dominate applications like foundational models and computer vision, the use of SSMs—especially Mamba—for neural operators in PDEs remains underexplored, and their theoretical connections and potential advantages are yet to be fully understood.
@@ 2 Related Work
Data-Driven PDEs. Recent advances in fluid dynamics and solving PDEs have led to architectures modelling continuous-time solutions and multiparticle dynamics[kochkov2021machine, lusch2018deep]. Physics-informed models now offer solutions in unsupervised and semi-supervised settings[raissi2019physics, li2020fourier]. These models typically encode spatial data and evolve over time, utilising methods like convolutional layers[ronneberger2015u, wiewel2019latent] symbolic neural networks[udrescu2020ai], and residual networks[he2016deep]. Finite element methods (FEM), including Galerkin and Ritz, are also integrated into learning frameworks[chen2021continuous].
@@
Neural Operators. Neural operators, such as the Graph Neural Operator[li2020neural] and Fourier Neural Operator[li2020fourier], excel at learning mappings in infinite-dimensional spaces, particularly by leveraging techniques like graph structures or transformations in Fourier space. The Fourier Neural Operator (FNO) and its variations, including the incremental, factorised, adaptive FNO, and FNO+[zhao2022incremental, tran2021factorized, guibas2021adaptive] have shown exceptional performance in both speed and accuracy. Their key advantage lies in their ability to maintain discretisation invariance, which sets them apart in many applications. DeepONet[lu2019deeponet] pioneered the nonlinear operator approximation using separate networks for inputs and query points, while extensions like MIONet handle multiple inputs[jin2022mionet]. Challenges like irregular grids are being addressed through grid mapping and subdomain partitioning[li2022irregular, wen2022u] though scalability for diverse inputs remains a key focus.
@@
Transfomers for PDEs. The Transformer model[vaswani2017attention] stands out due to its distinctive features, primarily its use of attention mechanisms to model the relationships among input elements. Initially, it was developed for NLP, and attention mechanisms have been adapted to PDEs, providing flexible and efficient mappings between function spaces. Recent theoretical work [yun2019transformers] also establishes that transformers are universal approximators of sequence-to-sequence functions, capable of representing arbitrarily complex dependencies. Galerkin attention[cao2021choose] introduced linear complexity to reduce computational costs, inspiring further developments like GNOT[hao2023gnot] and OFormer[li2022transformer], which achieve state-of-the-art results. Additionally, graph-based Transformers have also been explored to capture complex interactions in irregular domains[bryutkin2024hamlet].
@@
State-Space Models for PDEs & Comparison to Ours. Initial studies on SSMs for PDEs, like MemNO[buitrago2024benefits], explored combining FNO with S4 but were restricted to low-resolution or noisy inputs. In contrast, we introduce the Mamba Neural Operator, which generalises the SSM framework to neural operators, making it compatible with any architecture, including Transformers. Our approach extends the theoretical foundations for broad applicability to any PDE family, highlighting Mamba’s effectiveness in diverse scenarios. At the time of our submission, the work of that[hu2024state] proposed integrating state-space models into neural operators for dynamical systems. While related, our work differs significantly, their approach focuses on dynamical systems and tests only on ordinary differential equations (ODEs), whereas we target parametric partial differential equations (PDEs). Additionally, we provide a theoretical understanding showing that neural operator layers share a comparable structural framework with time-varying SSMs, demonstrating alignment between hidden space updates and the iterative process in neural operators.
3 Mamba Neural Operator
@@ 3.1 Problem Statement
3.2 Preliminaries: Transformer and Mamba
@@
of which the memory complexity is $O(n^{2})$. To reduce the computational inefficiency, Galerkin-type attention was proposed by[cao2021choose] to remove Softmax attention with linear complexity. It defines as follows:
$\operatorname{Attention}_{g}(Q,K,V)=\frac{Q(K_{g}^{T}V_{g})}{d},$ (4)
in which $K_{g}^{T}$ and $V_{g}$ denote layer normalisation of K and V, as described in[lei2016layer]. The Galerkin-type attention mechanism involves two matrix product operations, resulting in a computational complexity of $O(nd^{2})$. This reduces the sequence length dependency to only $O(n)$.
@@
where $A\in\mathbb{C}^{N\times N}$ and $B,C\in\mathbb{C}^{N}$ and $D\in\mathbb{C}^{N}$. Mamba, a more advanced variant of SSMs, refines this formulation by incorporating efficient state space parameterisation and selection mechanisms. Unlike earlier models such as S4, which uses bilinear method, Mamba adopts zero-order holds, allowing it to handle larger hidden states and longer sequences more effectively. This makes Mamba particularly well-suited for complex sequence modeling tasks, such as natural language processing and time-series analysis.
3.3 State Space Models Discretisation for PDEs
@@
We start by demonstrating that the discretisation of S6 (Mamba) is equivalent to the well-known Euler method when the Taylor series expansion is applied. Mamba utilises zero-order holds, resulting in the following discretisation, which reads:
@@
To discretise the continuous-time SSM, we apply the zero-order hold (ZOH) method, a standard control-theoretic discretisation technique that assumes the input signal $u(t)$ remains constant over each sampling interval. Given a continuous-time linear state-space system:
$\dot{h}(t)=Ah(t)+Bu(t),\quad y(t)=Ch(t),$ (7)
the ZOH discretisation produces an equivalent discrete-time system of the form:
$h_{k+1}=\tilde{A}h_{k}+\tilde{B}u_{k},$ (8)
where:
$\tilde{A}=e^{A\Delta},\quad\tilde{B}=\left(\int_{0}^{\Delta}e^{A\tau}d\tau\right)B=A^{-1}(e^{A\Delta}-I)B.$
This is the formulation used in Equation (6) of our paper. ZOH is particularly relevant to our work because it preserves the system dynamics while enabling a principled transition from continuous-time to discrete-time models—crucial when applying SSMs to neural operators operating on discretized PDE data. Under ZOH, the discrete-time equivalent of the system is given by:
$\tilde{A}=e^{A\Delta},\quad\tilde{B}=A^{-1}(e^{A\Delta}-I)B,$ (9)
where $\Delta$ is the step size. This formulation preserves the dynamics of the continuous system and forms the basis for our discrete SSM implementation in (6).
Discretisation of SSM. To incorporate the SSM into deep learning frameworks, we need to transform the continuous-time SSM into a discrete formulation. This is done by expressing the continuous-time system as an ordinary differential equation (ODE) and then solving it numerically. As discussed in[liu2024vmamba], the discrete SSM reads:
@@
$\begin{split}h(t+\Delta)&=\hat{A}h(t)+\hat{B}u(t)=(I+\Delta A)h(t)+\Delta Bu(t)\\ &=h(t)+\Delta(Ah(t)+Bu(t))=h(t)+\Delta h^{\prime}(t).\end{split}$ (12)
which implies it is a first order Euler method. It is straightforward to show that $\tilde{A}=\hat{A}$ since $\tilde{A}=e^{A\Delta}=I+A\Delta+O(\Delta^{2})=I+A\Delta=I+A\Delta=\hat{A}$. Similarly, we observe that $\tilde{B}=\bar{B}=\hat{B}$. Therefore, the discretisation used in the SSM method can be replaced with the zero-order hold method by substituting $\hat{A}=\tilde{A}$ and $\hat{B}=\tilde{B}$, we get:
$\begin{split}h(t+\Delta)&=\hat{A}h(t)+\hat{B}u(t)=\tilde{A}h(t)+\tilde{B}u(t)\\ &=(e^{\Delta A})h(t)+((\Delta A)^{-1}(e^{\Delta A}-I)\cdot\Delta B)u(t)\\ &=(I+\Delta A+O(\Delta^{2})h(t)+(\Delta B)u(t)\\ &=(I+\Delta A)h(t)+(\Delta B)u(t)\\ &=h(t)+\Delta Ah(t)+\Delta Bu(t)\\ &=h(t)+\Delta(Ah(t)+Bu(t))=h(t)+\Delta h^{\prime}(t).\end{split}$ (13)
@@
Why is Proposition 1 important for PDEs? Proposition 1, which establishes the equivalence between the Zero-Order Hold (ZOH) method and the Euler method, is crucial for understanding Mamba’s performance in solving partial differential equations (PDEs). This equivalence demonstrates that ZOH can be viewed as a more generalised and accurate variant of the Euler method. Throughout this work $\Delta>0$ denotes the time step for the mamba used to advance the hidden state of a continuous-time state–space model (SSM). Applying the forward-Euler scheme to the linear ODE $h^{\prime}(t)=Ah(t)+Bu(t)$ gives the following update
$h(t+\Delta)=h(t)+\Delta\bigl(Ah(t)+Bu(t)\bigr)$ (14)
, so Euler’s method uses a time increment $\Delta$. Proposition 1 shows that the Zero-Order Hold (ZOH) discretisation,
$\tilde{A}=\mathrm{e}^{A\Delta},\qquad\tilde{B}=\bigl(A^{-1}\!(\mathrm{e}^{A\Delta t}-I)\bigr)B,$
reduces to the Euler matrix pair $(I+\Delta A,\;\Delta B)$ when the matrix exponential is truncated at first order. Hence, ZOH can be viewed as a higher-order method of forward Euler: it retains the leading terms of the exact flow map $\mathrm{e}^{A\Delta}$ and therefore achieves a local error $O(\Delta^{2})$ instead of $O(\Delta)$. In the supplement, we provided an example to explain the difference between the Euler method and the ZOH method.
In this work, we provide the first analysis of the Zero-Order Hold method, showing that it could be extended as a higher-order method via Taylor series expansion and reduces to the first-order Euler scheme when truncated at first order. To the best of our knowledge, this perspective is the first work that provides a connection between classical numerical methods and Zero Order Hold method. Although the Forward Euler method is the simplest first-order scheme, it is only conditionally stable and limited in accuracy. Higher-order methods can achieve faster convergence and improved accuracy, though often with different stability characteristics. By reforming ZOH as a Taylor series method, we open the possibility to extend it to higher-order methods. This is particularly important because the choice of time step and ODE solver directly governs how the state-space model propagates forward in time, and thus has a critical impact on the accuracy of PDE solutions.
3.4 Network Architecture
@@
As depicted in Figure 1, the data processing pipeline in our Mamba Neural Operator (MNO) is composed of three key stages: Bi-Directional Scan Expand, S6/Cross S6 Block, and Bi-Directional Scan Merge. As shown in Figure 2, our method uses Mamba and bidirectional scan to achieve linearly complex and global receptive fields, effectively combining the strengths of CNNs and Vision Transformers (ViTs). Unlike CNNs, which capture only local dependencies, and ViTs, which model global context but come with quadratic complexity, Mamba (2D) introduces a bi-scan mechanism that enables efficient global information propagation. This design provides linear complexity to CNNs while retaining the expressive global modeling ability of ViTs. When solving PDEs over a fixed grid, the input data can be structured as grid-based data, similar to an image. In the first stage, Bi-Directional Scan Expand, the MNO unfolds the input data into sequences by traversing the grid along two distinct paths. These sequences, representing input patches, are processed independently in the next step. The second stage, S6/Cross S6 Block, involves processing each patch sequence using either an S6 or Cross S6 block, depending on the model variation being employed. For instance, in the enhanced version of Mamba, the GNOT model utilises a Cross S6 block followed by an S6 block for further refinement. Finally, in the Bi-Directional Scan Merge stage, the processed sequences are reshaped and merged back together to generate the output map, completing the data forwarding process. This structured approach allows the MNO to efficiently handle grid-based input data, enabling scalable solutions for PDEs.
@@
[…]
@@
3.5 Mamba for Neural Operators
Neural Operators[li2020neural] aim to learn mappings between function spaces, providing a framework for solving partial differential equations (PDEs) and other problems involving continuous functions. It updates the value by an iterative method: $i_{0}\rightarrow i_{1}\rightarrow\ldots\rightarrow i_{T},$ where each $i_{j}$ (for j=0,1,,T1j=0,1,...,T-1) maps to $\mathbb{R}^{d_{v}}$. Let the input be $a(x)$ and the output be $u(x)$ . The input $a$, drawn from set $A$, is initially lifted to a higher-dimensional representation: $v_{0}(x)=P(a(x))$ where $P$ is a local transformation, typically parameterised by a fully-connected neural network. We then apply iterations to update $i_{t}\rightarrow i_{t+1}$ as defined in Definition 1. The final output: $u(x)=Q(v_{T}(x))$ is the result of projecting $v_{T}$ via the transformation: $Q:\mathbb{R}^{d_{v}}\to\mathbb{R}^{d_{u}}.$ Each update from $i_{t}$ to $i_{t+1}$ involves the integration of a non-local integral operator $K$ and a local nonlinear activation function $\sigma$. One of the main results of this work is establishing the equivalence between neural operators and the Mamba framework. Therefore, we first introduce fundamental definitions stated in [li2020neural] that are essential for demonstrating this relationship.
Figure 1: (A) Illustration of Mamba Neural Operator. Input image patches are processed by following two distinct scanning paths (referred to as Bidirectional -Scan). Each sequence generated from these paths is passed through separate S6 blocks/ Cross S6 Blocks for independent processing. Afterwards, the outputs from the S6 blocks / Cross S6 Blocks are combined to form a feature map, resulting in the final output (Bidirectional-Merge). (B) and (C) are the detailed blocks of the S6 Block and Cross S6 Block, respectively.